OpenStudy (zubhanwc3):

The differential equation dy/dx = (x+1)/(x-2)

OpenStudy (zubhanwc3):

The differential equation \[\frac{ dy }{ dx } = \frac{ x+1 }{ x-2 }\] I. produces a slope field with horizontal tangents at x = -1 II. produces a slope field with vertical tangents at x = 2 III. produces a slope field with rows of parallel segments I only II only I and II only all of them II and III only

OpenStudy (tkhunny):

I. No. It's at x = +1 II. Yes. x = 2 makes the denominator explode. III. You tell me.

OpenStudy (zubhanwc3):

i would assume that III is correct, because is vertical tangents, wouldnt it be parallel?

OpenStudy (tkhunny):

?? Not sure that makes any sense.

OpenStudy (zubhanwc3):

actually mb, its not parallel because tangents arent straight lines are they?

OpenStudy (tkhunny):

?? You seem to be struggling with exactly what a Slope Field is. As we get away from x = 0, far away, the slope tends toward +1 At x = -1, we have 0/(-3) = 0. So actually, I is good. It is tangents are horizontal for x= -1. You should have argued with me on that point.